During the course of a​ year, the labor force consists of the same 963 people. Of​ these, 15 lack skills that employers desire and hence remain unemployed throughout the year. At the same​ time, every month during the​ year, 39 different people become​ unemployed, and 39 other different people who were unemployed find jobs. There is no cyclical unemployment.
Answer #1

Solution:-

Calculate the frictional unemployment rate -

Frictional unemployment = 39

labor force = 963

Frictional unemployment rate = (Frictional unemployment/Labor force)*100

                                           = (39 / 963) * 100

                                           = 4.05%

The frictional unemployment rate is 4.05%.

Calculate the unemployment rate -

Total unemployment = Frictional unemployment + Structural unemployment

                              = 39 + 15

                              = 54

Unemployment rate = (Total unemployment/Labor force) * 100

                             = (54 / 963) * 100

                             = 5.61%

The unemployment rate is 5.61%.